OpenStudy (anonymous):

Solve each system of equations... 4w-v=15 -2w+6v=-2

OpenStudy (luigi0210):

First what you would do is multiply the bottom equation by 2, giving you: 4w-v=15 -4w+12v=-4 Now add the two equations together.. the w's would cancel: 11v=11 now divide v=1 Now plug in v and solve for w 4w-1=15 (add one) 4w=16 (divide) w=4

OpenStudy (anonymous):

Thank you so much I have been trying to get help for at least a week now

OpenStudy (luigi0210):

Your welcome, do you understand what to do?

OpenStudy (anonymous):

no it makes no sense to me at all.

OpenStudy (luigi0210):

Well there are 3 ways to solve these equations: By canceling, substitution, and graphing

OpenStudy (anonymous):

It still makes no sense to me either way you are like the fifth person to try and explain this to me and it makes no sense

OpenStudy (luigi0210):

Is there any way I can make it simpler for you?

OpenStudy (anonymous):

can you help me with this one? -5m+3n=12 5m-8n=23

OpenStudy (anonymous):

you can try actually but I don't know if it will help

OpenStudy (luigi0210):

What is one thing you could do? Can you cancel anything out?

OpenStudy (anonymous):

the -5 and 5

OpenStudy (luigi0210):

Yes, so what does that leave us with when you add the two equations?

OpenStudy (anonymous):

would you add or subtract it?

OpenStudy (luigi0210):

You always add the equations

OpenStudy (luigi0210):

so what's 3n+(-8n)? and 23+12?

OpenStudy (anonymous):

ok so I got 11m=35

OpenStudy (luigi0210):

That's a -8n.. it's the same as 3n-8n.. so it would actually be -5n=35 Be careful with the positive and negative signs

OpenStudy (anonymous):

yeah I figured that out

OpenStudy (luigi0210):

Okay, so now divide both sides by -5 to get the variable alone

OpenStudy (anonymous):

ok so I got -7 I plugged that in and got 21n=12 and 56n=23

OpenStudy (luigi0210):

Yes, but all you have to do is just plug it in to just any one equation

OpenStudy (anonymous):

ok so I still don't get the problem

OpenStudy (luigi0210):

-5m+3n=12 Plug in -7 into this equation as a substitution for n -5m-3(7)=12 -5m-21=12 -5m=32 m=-5/32

OpenStudy (luigi0210):

my mistake that's suppose to be 33 not 32, sorry

OpenStudy (anonymous):

ok wanna help me with another one haha

OpenStudy (luigi0210):

sure

OpenStudy (anonymous):

ok 7j-3k=-14 2j+3k=32

OpenStudy (anonymous):

ok I think I get this one Im just doing these numbers in a random order lol

OpenStudy (luigi0210):

Alright, so what did you get?

OpenStudy (anonymous):

ok so I just got j=-2 and j=16

OpenStudy (luigi0210):

why did you get 2 answers for j?

OpenStudy (anonymous):

idk

OpenStudy (luigi0210):

I'm sorry, I can't really help anymore..

OpenStudy (anonymous):

ok